OpenStudy (anonymous):

Find the slope of the line that passes through (-9, 5) and (6, 13)

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

k :3 thank you :3 wait stay here so i can ask if its wrong or something when im finished :0

hartnn (hartnn):

sure!

OpenStudy (anonymous):

k thanks hold on :3

OpenStudy (anonymous):

i got 8/-3 did i do it wrong?

hartnn (hartnn):

the denominator , 6 - (-3) = ... ? numerator =8 is correct :)

OpenStudy (anonymous):

but the numerator is-9?

hartnn (hartnn):

***6- (-9) = yeah sorry, typo

hartnn (hartnn):

numerator is not -9, numerator = y1-y2 = 13-5 = 8 is correct.

OpenStudy (anonymous):

oh okay yeah so 6-(-9) would be -3... no?

OpenStudy (anonymous):

or just 3?

hartnn (hartnn):

nopes, 6- (-9) = 6+9 = .. two negatives become a positive!

OpenStudy (primeralph):